OpenStudy (anonymous):

log3( log2 (logx 5))=0 ??

OpenStudy (shubhamsrg):

make use of these : 1) if log a b =c ,, (where a is base) then a^c =b 2) log 1 =0 (for any base) this should help..

OpenStudy (zehanz):

Try to do this step by step: the last calculation yields 0: \[\log_{3}(...)=0 \]so according to @shubhamsrg 's tip, the dots (...) are equal to 1. This means:\[\log_{2}(...)=1\] Now the dots (...) are equal to 2^1 = 2. See? Maybe you can try the last step yourself!

OpenStudy (zehanz):

OK, now we have:\[\log_{x} 5=2\]which means x² = 5, so x = ...

OpenStudy (anonymous):

Thanks, I get it now.

OpenStudy (anonymous):

aliter log3( log2 (logx 5))=0 [bcoz log 1=0] hence ( log2 (logx 5))=1 [bcoz log2 2=1 read as log of 2 base 2] hence logx 5=2 or x^2 = 5 or x= sqrt(5)

OpenStudy (anonymous):

thank you